OpenStudy (anonymous):

I'm finding arc lengths. y=x^(2/3) and coordinates are (8sqrt3/9, 4/3) and (0,0)

OpenStudy (anonymous):

\[y \prime= \frac{ 2 }{ 3x ^{1/3}}\]

OpenStudy (anonymous):

do you know the anwser to this question what is the 6th term of (2x-3y)^11?

OpenStudy (anonymous):

\[(y \prime)^{2}= \frac{ 4 }{ 9x ^{2/3} }\]

OpenStudy (anonymous):

the 6th term?

OpenStudy (anonymous):

a.7,187,024x^6y^5 B.-7,185,o24x^6y^5 C.7,185,024x^6y^5 D.-7,185,024x^6y^5

zepdrix (zepdrix):

hi c:

OpenStudy (anonymous):

\[\sqrt{\frac{ 9x ^{2/3}+4 }{ 9x ^{2/3} }}\]

OpenStudy (anonymous):

Hey! Someone was trying to jack my question, haha

zepdrix (zepdrix):

ya i see that -_-

zepdrix (zepdrix):

So to find arc length ...

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

we break up our curve into little pieces.